OpenStudy (anonymous):

LOST Identify the real and imaginary parts, respectively, of a complex number. 5 - (-2i) Select one: a. -5 and -2 b. 5 and 2 c. -2 and 5 d. 2 and -5

OpenStudy (anonymous):

i am going to bet that this is really \[5+2i\] making the real part 5 and the imaginary part 2

HanAkoSolo (jamierox4ev3r):

think of it as distributing the negative sign.

OpenStudy (anonymous):

and i don't believe you are lost, since the only trick here was that \[5-(-2i)=5+2i\]

OpenStudy (anonymous):

I dont get how to solve that or know how you even began to set that up

OpenStudy (mathmale):

If you're given a complex number, the real part is the part that does not have the operator i associated with it. The imaginary part does have the operator i associated with it. These are definitions that need to be learned and remembered.

OpenStudy (anonymous):

If its 3+6i = 3+ (-6)??

HanAkoSolo (jamierox4ev3r):

not necessarily

OpenStudy (anonymous):

Then i dont get how to do that

HanAkoSolo (jamierox4ev3r):

3+ (-6) is equal to 3-6

OpenStudy (anonymous):

oh dear there is nothing to set up \[5-(-2x)=5+2x\] right ?

OpenStudy (anonymous):

i.e. \[-(-2)=2\]

OpenStudy (mathmale):

Starting with 5 - (-2i): Please focus on the -(-2i). This simplifies to what? What is (-1)(-1)? I agree with satellite that 5 - (-2i) can be re-written as 5 + 2i. This complex number has two terms. Which term is real (has no i associated with it)?: Which term is imaginary (has i associated with it)?:

OpenStudy (mathmale):

If you can answer those two questions, you'll be done.

OpenStudy (anonymous):

if there both positive it stays the same right?

HanAkoSolo (jamierox4ev3r):

remember these ground rules:, **a positive number subtracted by a negative number will ALWAYS be positive Example : 5-(-6)=11 **a negative number subtracted by a positive number will ALWAYS be negative Example: -5-(6)=-11 **a positive number subtracted by a positive number...is just basic math really Example: 5-6=-1 (since the 6 is greater than 5, the answer is negative) **a negative number subtracted by a positive number...this one is tricky! Example:-5-(-6)=1 remember, when there are two negative signs next to each other, the operation becomes like addition...so this could be rewritten as -5+6, which equals 1

HanAkoSolo (jamierox4ev3r):

^^hope this clears things up

OpenStudy (anonymous):

what about a positive and positive theyll stay the same right?

HanAkoSolo (jamierox4ev3r):

you mean if a positive is subtracted by a positive?

OpenStudy (anonymous):

no its added??? 3+6i

HanAkoSolo (jamierox4ev3r):

well then yes, it will always remain positive.

HanAkoSolo (jamierox4ev3r):

addition is pretty straight forward, but I could break that down for you too ^^ I only broke down subtraction up there for you lol

OpenStudy (anonymous):

for this one its be D because both would be 0 -4 - (- 4i) Select one: a. -4 and 4 b. 4 and 0 c. 0 and 4 d. None of the above

OpenStudy (mathmale):

Starting with 5 - (-2i): Please focus on the -(-2i). This simplifies to what? What is (-1)(-1)? I agree with satellite that 5 - (-2i) can be re-written as 5 + 2i. This complex number has two terms. Which term is real (has no i associated with it)?: Which term is imaginary (has i associated with it)?:

HanAkoSolo (jamierox4ev3r):

More ground rules: **A positive number added to a positive number is ALWAYS positive Example: 5+6=11 **A positive number added to a negative number...depends Example:5+(-6)=-1 this can be rewritten as 5-6, which equals -1 **A negative number added to a positive number...depends as well Example: -5+6=1 Since 6 is large enough for -5 to become positive, 1 is the answer **A negative number added to a negative number is ALWAYS negative Example -5+(-6)=-11 this can be rewritten as -5-6, which is equal to -11

OpenStudy (anonymous):

for this one its be D because it would equal 8 but thats not an answer?? -4 - (- 4i) Select one: a. -4 and 4 b. 4 and 0 c. 0 and 4 d. None of the above

HanAkoSolo (jamierox4ev3r):

Are you still trying to identify the imaginary/real parts of the number?

OpenStudy (anonymous):

yes

HanAkoSolo (jamierox4ev3r):

In that case, I'm sorry to say that you may want to think again. Think in this manner: 

OpenStudy (anonymous):

oh so its 4 and -4

HanAkoSolo (jamierox4ev3r):